# EfficientZero V2: Mastering Discrete and Continuous Control with Limited Data

### v1

@@ Appendix F Proof for SVE Error Bound
To provide detailed proof for the upper bound of the MSE of the search-based value estimation $\hat{V}_{\text{S}}$, we follow a structured approach inspired by Model-based Value Estimation (MVE) (Feinberg et al., 2018), with adjustments for the specifics of MCTS.
@@
According to $L-Lipschitz$ continuity of $\mathcal{G}$ and $\mathcal{R}$, we have
$\mathbb{E}\left[(\hat{r}_{t}-r_{t})^{2}\right]$ $\leq L_{r}^{2}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]$
$\mathbb{E}\left[\left(\hat{V}(\hat{s}_{t})-V_{\pi}(s_{t})\right)^{2}\right]$ $\leq L_{V}^{2}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]$
@@
$\mathbb{E}\left[(\hat{r}_{t+h}-r_{t+h})^{2}\right]$ $\leq L_{r}^{2}\mathbb{E}\left[\|\hat{s}_{t+h}-s_{t+h}\|^{2}\right]$
$\mathbb{E}\left[\left(\hat{V}(\hat{s}_{t+h})-V_{\pi}(s_{t+h})\right)^{2}\right]$ $\leq L_{V}^{2}\mathbb{E}\left[\|\hat{s}_{t+h}-s_{t+h}\|^{2}\right]$
Using Cauchy-Schewartz inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$, we have
$\begin{split}\text{MSE}_{\nu}(\hat{V}_{\text{S}})\leq&\frac{2}{N^{2}}\sum_{n=0}^{N}\Bigg{(}\sum_{t=0}^{H(n)}\gamma^{2t}(\hat{r}_{t}-r_{t})^{2}+\gamma^{2H(n)}\left(\hat{V}(\hat{s}_{H(n)})-V_{\pi}(s_{H(n)})\right)^{2}\Bigg{)}\\ \leq&\frac{2\epsilon^{2}}{N^{2}}\sum_{n=0}^{N}\left(\sum_{t=0}^{H(n)}\gamma^{2t}L_{r}^{2}+\gamma^{2H(n)}L_{V}^{2}\right)\end{split}$ (24)
Considering whether the coefficient part
$\frac{2}{N^{2}}\sum_{n=0}^{N}\left(\sum_{t=0}^{H(n)}\gamma^{2t}L_{r}^{2}+\gamma^{2H(n)}L_{V}^{2}\right)$ (25)
converges with increasing search depth $H(n)$, we separate it into two parts:
@@
    The reward error term $\sum_{t=0}^{H(n)}\gamma^{2t}L_{r}^{2}$: Given $L_{r}\in c$ and $\gamma\in(0,1)$, the series is obviously convergent.
@@
    The terminal state value error term $\gamma^{2H(n)}L_{V}^{2}$ is similarly converged to 0 if $H(n)\to\infty$.
On the other hand, this upper bound will also converge to 0 if the model is optimal $\epsilon\to 0$. ∎
@@ I.2.6 BBF
Additionally, we include an ablation study focusing on the sampling-based Gumbel search and the mixed value target. Figures 9 and 10 demonstrate that our search method and mixed value target achieve superior performance on tasks with proprioceptive and image inputs. Notably, our search method maintains strong performance in relatively easy tasks even with limited search simulations.
@@
Figure 9: The ablation study of our search method, namely sampling-based Gumbel search (S-Gumbel Search). Besides, we compare with sample MCTS (Hubert et al., 2021), and our search method with different simulations.
Figure 10: The ablation study of the mixed value target we propose. Besides, we compare with multi-step TD target and double value target. Double Q-value target is similar to the value target in DDPG (Lillicrap et al., 2015) (more details can be found in Section 5.3).

### v2

@@ Appendix E Proof for SVE Error Bound
To provide detailed proof for the upper bound of the MSE of the search-based value estimation $\hat{V}_{\text{S}}$, we follow a structured approach inspired by Model-based Value Estimation (MVE) (Feinberg et al., 2018), with adjustments for the specifics of MCTS and considerations of errors of reward and value estimations.
@@
According to $L-Lipschitz$ continuity of $\mathcal{R}$ and Cauchy inequality, we can derive
$\mathbb{E}\left[(\hat{r}_{t}-r_{t})^{2}\right]$ $=\mathbb{E}\left[(\mathcal{R}(\hat{s}_{t})-\mathcal{R}^{*}(s_{t}))^{2}\right]$
$=\mathbb{E}\left[(\mathcal{R}(\hat{s}_{t})-\mathcal{R}(s_{t})+\mathcal{R}(s_{t})-\mathcal{R}^{*}(s_{t}))^{2}\right]$
$\leq 2\mathbb{E}\left[(\mathcal{R}(\hat{s}_{t})-\mathcal{R}(s_{t}))^{2}\right]+2\mathbb{E}\left[(\mathcal{R}(s_{t})-\mathcal{R}^{*}(s_{t}))^{2}\right](\text{Cauchy inequality})$
$\leq 2L_{r}^{2}\mathbb{E}\left[\|\hat{s}_{t}-s_{t}\|^{2}\right]+2\mathbb{E}\left[\|\mathcal{R}(s_{t})-\mathcal{R}^{*}(s_{t})\|^{2}\right]$
$\leq 2L_{r}^{2}\epsilon_{s}^{2}+2\epsilon_{r}^{2}$
Similarly, we derive the error bound of value estimation as
$\mathbb{E}\left[\left(\hat{V}(\hat{s}_{t})-V_{\pi}(s_{t})\right)^{2}\right]=\mathbb{E}\left[(\mathcal{V}(\hat{s}_{t})-\mathcal{V}^{*}(s_{t}))^{2}\right]\leq 2L_{v}^{2}\epsilon_{s}^{2}+2\epsilon_{v}^{2}$
@@
$\begin{split}\text{MSE}_{\nu}(\hat{V}_{\text{S}})\leq&\frac{2}{N^{2}}\sum_{n=0}^{N}\Bigg{(}\sum_{t=0}^{H(n)}\gamma^{2t}(\hat{r}_{t}-r_{t})^{2}+\gamma^{2H(n)}\left(\hat{V}(\hat{s}_{H(n)})-V_{\pi}(s_{H(n)})\right)^{2}\Bigg{)}\\ \leq&\frac{4}{N^{2}}\sum_{n=0}^{N}\left(\sum_{t=0}^{H(n)}\gamma^{2t}(L_{r}^{2}\epsilon_{s}^{2}+\epsilon_{r}^{2})+\gamma^{2H(n)}(L_{V}^{2}\epsilon_{s}^{2}+\epsilon_{v}^{2})\right)\end{split}$ (28)
Considering the convergence with increasing search depth $H(n)$, we separate it into two parts:
@@
    The reward error term $\sum_{t=0}^{H(n)}\gamma^{2t}(L_{r}^{2}\epsilon_{s}^{2}+\epsilon_{r}^{2})$: Given $L_{r}\in\mathcal{C}$, $\gamma\in(0,1)$, and $\epsilon_{s},\epsilon_{r}$ decaying with training, the series is obviously convergent.
@@
    The terminal state value error term $\gamma^{2H(n)}(L_{V}^{2}\epsilon_{s}^{2}+\epsilon_{v}^{2})$ is similarly converged.
On the other hand, this upper bound will also converge to 0 if the model is optimal $\epsilon_{s},\epsilon_{r},\epsilon_{v}\to 0$. ∎
@@
Appendix F Details of Mixed Value Target
The mixed value target is calculated for each transition sampled from the buffer. It contains two types of value targets: the search-based value target and the multi-step TD target.
Specifically, we use two criteria to determine if we should use a search-based value target. The first criterion is whether the sampled transition comes from recent rollouts. If the transition is not from recent rollouts, meaning the transition is considered stale, we opt for the search-based value as the target. Otherwise, the multi-step TD (Temporal Difference) target is employed. The second criterion considers whether the current training step is in the early stage. During this phase, all transitions in the buffer are fresh. Additionally, the error in the dynamic model is still significant, resulting in inaccuracies in search-based value estimation. Therefore, we opt for the multi-step TD target as the target value.
@@ I.2.6 BBF
J.2 Ablation
Additionally, we include an ablation study focusing on the sampling-based Gumbel search and the mixed value target. Table 4, 5 and 6 demonstrate that our search method and mixed value target achieve superior performance on tasks with proprioceptive and image inputs. The action space is 1-dimensional in tasks such as Acrobot Swingup, Cartpole Swingup Sparse, and Pendulum Swingup. The dimension is greater than 2 in other DM-Control tasks. For Atari 100k, we selected 3 of the relatively most challenging tasks for our ablation study.
We have observed that our mixed value target outperforms other value estimation methods in most tasks, such as multi-step value target and generalized advantage estimation (GAE). This indicates that the method mitigating the off-policy issue consistently performs better. Compared to Sample MCTS, the S-Gumbel Search method significantly enhances performance in tasks with a high-dimensional action space. The results demonstrate that the S-Gumbel Search method can achieve superior performance with the limited simulations.
Table 4: Proprio Control 50-100k
Method Acrobot swingup
Cartpole
swingup sparse
Pendulum swingup Reacher hard Walker walk Walker run Quadruped walk
Original EZ-v2 297.7 795.4 825.4 795.4 944.0 657.2 925.8
W/ Multi-step TD Target 256.7 473.6 836 590.7 839.3 521.6 649.7
W/ GAE 275.1 766.7 336 601.6 757.3 512.7 719.7
W/ Sample MCTS 248.1 789.3 357.4 754.7 691.7 381.1 254.7
Table 5: Vision Control 100-200k
Method Acrobot swingup
Cartpole
swingup sparse
Pendulum swingup Reacher hard Walker walk Walker run Quadruped walk
Original EZ-v2 231.8 763.6 726.7 961.5 888.8 475.3 433.3
W/ Multi-step TD Target 186.3 631.7 411.4 878 880.5 496.7 410.0
W/ GAE 122.0 796.1 718.0 934.9 765.1 491.5 299.3
W/ Sample MCTS 73.9 786.1 372.9 938.8 619.9 264.2 141.4
Table 6: Atari 100k
Game
Method Asterix UpNDown Qbert
Original EZ-v2 61810.0 15224.3 16058.3
W/ Multi-step TD Target 26023.1 13725.7 9463.3
W/ GAE 22816.7 16255.3 7807.3
J.3 Computation Load
We have included practical comparisons between the TD-MPC2 and EZ-V2 algorithms in terms of computational load, using the ’Walker Run’ task as an example. The following results include the parameter count, FLOPs (per decision step) and training time for each algorithm. The methods were benchmarked on a server equipped with 8 RTX 3090 graphics cards.
Table 7: Model Parameters and Training Performance
Parameters Flops per decision step Time per 100k training (h)
EZ-V2 $\mathbf{1.3\times 10^{6}}$ $\mathbf{4.7\times 10^{7}}$ 2.7
TD-MPC2 $4.9\times 10^{6}$ $3.6\times 10^{10}$ 3.3
EZ-V2 requires 1000 times fewer FLOPs per decision step compared to TD-MPC2, while also using almost 4 times fewer parameters. The decision step, crucial for collecting interaction data, occurs during the evaluation. This remarkable efficiency arises from two primary factors:
  • The planning process in TD-MPC2, utilizing the MPPI method, involves predicting 9216 latent states. In contrast, our method extends only 32 latent states, significantly reducing the computational load.
  • TD-MPC2 relies on an ensemble of Q-functions (5 heads), and its latent state dimension is 4 times larger than our method’s, leading to higher computational demands.
The FLOPs per decision step are vital for deployments, especially in robotics control. Since robots typically have limited edge computing resources, using less computation is essential for real-time tasks. TD-MPC2 needs more computational resources and high-performance computing. In contrast, our method runs efficiently with less demand of computing resources, making it ideal for real-time robotic planning.
Regarding training time, both methods require a similar amount of time per 100k training steps. This similarity in time consumption is due to that EZ-V2 and TD-MPC2 share similar unrolled training frameworks with the same batch size. EZ-V2 wins in speed due to its distributed implementation, without rollout and evaluation overheads in the training process.